PartnerObserving Log for 2007 Jan 23/24

Observers: Rick Pogge
Inst. Support: Olga Kuhn
TOs: Aaron Ceranski

John Hill is also present tonight.

UTC Jan 23 2100: At Ray Bertram's direction, Aaron began filling the dewar at about 2030UT (Jan23). Temperature began to drop and hit 170K about 30m later. Pressure dropped during the same interval, and is now around 5e-5 mbar and falling steadily.

All times below are UTC 2007 Jan 24

0000 - Cryostat P & T now at145K and 1.2e-5, with a flat enough slope to proceed

0025 - Fired up LBC-B w/o problems, running the 25bias.ob sequence. Biases look normal, but only 22/25 made it into the repository.

0040 - Taking dark frame with dark900.ob in a darkened dome.

0043 - Sunset, sky is clear, humidity 51%, wind 9m/s out of NE, temp -6C, finally...

0057 - Opening the dome...

0100 - Sky flat U test, pointing to RA=02:10, Dec=+27 - test, 6x, then SkyFlatU _5.ob, 6s dithering started at 0115 (exec time ~7min)

0123 - B Sky Flat 5s test image then 5x10s dithering

0131 - V Sky Flat 10s test then 5x30s dithering (started 0134)

0145 - Going to Ryan 15-deg jan epoch 1 field (15deg_jan_epoch1.ob) - because it's the only thing up in the queue at this time.

Cripes. First time ever observing with the LBT and I'm using it to look for rocks...

0204 - started 15deg_jan_epoch1.ob, WFS setup took a bit more tweaking, dialing in Z11 with John Hill calling the numbers. Seeing ~1.25-arcsec on the first image in the sequence. Sky level ~14000ADU in V.

0215 - 10deg_jan_epoch1.ob started, little elevation change so now WFS check. Seeing ~1.5 arcsec

0229 - 5deg_jan_epoch1.ob after a WFS check, started at 0236... Mirror still stuck on afternoon temp, waiting for software guy to call back on the thermal computer is unresponsive. Maybe seeing will get better once it's back online... Done at

0247 - LBT going to zenith so John can reboot the thermal computer.

0255 - Standard star field L95 U, BV, & gr after WFS correct U start at 0307 (long slew because at zenith for thermal computer reboot). Mirror currently too warm by +4°C, so it'll be a while before things settle down. Seeing about 1.5 arcsec at V

0322 - Standard star field PG0231, WFS correct then U @ 0327, BV @ 0331 & gr @ 0335. Primary now +3°C too warm, but getting better

0339 - Standard star field SA98, WFS correct then U @ 0349, BV @ 0354 , & gr @ 0357 (yeah, I know, it was supposed to be SA92, that menu for selecting OBs show so much other stuff the mistake was easy in the very crowded file select dialog box, and I'd already done the WFS correction).

0403 - Ryan 15-deg field epoch 2 a little over 1h later as per the program. WFS correction then start @0407, then started 15-deg epoch 2 OB at 0412. Seeing still 1.5 arcsec. Mirror now +2°C and continuing to improve.

0424 - Ryan 10-deg field epoch 2. WFS correction start @0425, then started 10-deg epoch OB at 0434. But images are slightly trailed on the 2nd of the two V images, the first looks fine. John Hill says it's a rotator problem.

NOTE TO CHRIS KOCHANEK - the moon avoidance flag in the queue scheduler seems to be overly aggressive and is making selecting programs difficult because the word "moon" obliterates the airmass entry, even though the moon is at low phase and pretty far away.

0449 - Ryan 5-deg field epoch 2 - seeing is still in excess of 1.5 arcsec, so choosing targets with tighter seeing restrictions is why I'm doing more of this. Starting at 0454 after WFS correction. First image badly trailed. Second image in sequence aborted at 0502. John Hill is consulting the logs, and the rotator things it's happy, not clear what the problem is. Images taken at similar high elevation angle (>80\xB0) were fine (see above).

0503 - Moving to Sigma Ori Field 1 (for real this time) since it can handle 1.5 arcsec seeing while John puzzles about the rotator. WFS correction a little slow because it is taking ~30sec or more sometimes for the images to get moved from the archive to the respository for analysis. The logs show the transit from the instrument to the Archive is only a few seconds.

0512 - started Sigma Ori Field 1 OB 1 of 4. Getting a very bright glint on chip 1 from Sigma Ori in the U images, but otherwise images are 1.5+ arcsec and around (no rotator problems at this El=53\xB0). There is also extra light from Sigma Ori on chip 4 appearing as a broad fan of light adding about 700ADU per pixel over and above the typical U sky level on the images of around 1500ADU.

sigorif1_straylight.jpg
Sample mosaic showing the effect of light from Sigma Ori on offset field #1 in U.

0534 - started Sigma Ori Field 1 OB 2 of 4. No WFS tweak since we're at the same elevation (about 54\xB0). The wind has picked up some since sunset, now around 10-13m/s, and the humidity is up around 60%. Pattern of the glint from Sigma Ori on chip 2 changes with the dither pattern and filter, though the artifact in V are less dramatic looking thanin U. These data should be reviewed before more fields in this program are acquired to make sure the science goals aren't compromised by these artifacts. I'll leave that to my off-site secondary observer (Xiaohui or Jill) to assess tomorrow morning.

0606 - am only executing the first 2 images of the first of the 90-degree rotated images of Sigma Ori Field 1 in U (first 2 images of sdt_muzerolle_3.ob), and will terminate so we can assess how the light from Sigma Ori affects the rotated field. I am reluctant to spend more time on this field beyond the first 2 OBs given the background artifacts from the bright star. OB 3 of 4 (sdt_muzerolle_3.ob) was paused after the second image in the sequence at 90-degree field rotation at 0613.

0615 - Going to Skillman HOII, since that can tolerate >1-arcsec seeing, and we're about 1.25 arcsec. Aaron had to restart the azimuth drive when it began to oscillate. Went to the target w/o problems after the restart. Starting with WFS correction @ 0618. Mirror temperature is now <1°C above ambient, so we're almost there. Given the wind coming up, this is likely the least of our worries.

0627 - WFS took 3 iterations, started HOII_U1.ob. ... not getting any guide stars. Huh? Pausing after the first 300s exposure to assess... Z is way out of wack and the mirror is way out of focus. Correcting... Taking another 300s exposure. Pilot Error (I may have started the OB before the WFS correction was uploaded - trying again).

0636 - Restarted, looks much better, for dither 2 with the adjusted mirror focus. I am letting it execute dithers 3 and 4, and then I will go back and repeat dither 1 of 4 to complete the OB before proceeding with H0II_U2.ob Seeing, however, is degrading rapidly, now probably 2-arcsec.

0701 - started HOII_U2.ob for the 2nd half of this observation. Hmm, no guide star? Stopped. Reloaded OB and restarted at 0703. There is a star on the technical chip near (but not too near) one edge, but it does not lock onto it.

0709 - Doing WFS correction - not all that big of a focus change. We just can't seem to lock on a guide star. The one we got before must have been marginal, and subsequent images show the star moving more off the technical chip, which we don't understand. We tried editing the OB to introduce an offset of about 20arcsec, first guess was the wrong way. Second guess didn't help either. We're stumped - this makes no sense to us. Hacking OBs isn't a good idea, so we're moving on since we cannot find a way to get it to guide. The seeing is now too bad to continue on this target, and we probably didn't get good data for the first OB given the deteriorating seeing.

0731 - getting 3rd epoch (a wee late) of Ryan 15-deg field. Seeing now ~2 arcsec, and this is one of the few programs that will tolerate it. And we can finish a set of data, unlike our other programs tonight thus far.

0746 - setup and under way. Seeing is poor, so the WFS tweak up require extra iterations (oscillating spherical terms in first couple) before it stabilized.

0801 - Ryan 10-deg field, 3rd epoch, WFS tweak up, then started at 0808. Wind is really picking up, gusts now up to 20m/s. Aaron going to check the wind speed.

0825 Wind over limits, measured gusts to 55 mph, and some blowing snow so we're closing.

Never re-opened. Wind is not getting better and snow is definitely blowing around.

Night Summary

Conditions were clear and photometric all night. Winds increased steadily from sunset until we were eventually shutdown by high winds. Seeing started out at around 1.2-arcsec and then steadily degraded throught the night, eventually >2arcsec before we shut down.

Poor seeing conditions led me to do those programs that were most tolerant of >1.2 arcsec seeing.

  • Did Ryan program, 15-degree and 10-degree fields for 3 epochs. Attempt to do the 5-degree field was compromised by rotator problems during the 2nd epoch and we did not attempt a 3rd epoch (given wind problems, we would not have been able to if we wanted). I have indicated the 15-deg and 10-deg fields as "done" in the queue schedule.

  • Attempted Sigma Ori Field 1, but we encountered problems with scattered light from Sigma Ori proper (the fields are offset to keep this bright star off the field). Completed the first half of the 4 OBs for Field 1 (OB 1 and 2), and only took two images from OB3 (first of the 90-degree rotated fields) to show how the scattered light pattern across the chips changed with rotator angle. We will need to review this program to see what modifications are needed.

  • Attempted Skillman HOII field. Two problems emerged. One was that there is a serious lack of guide stars on the technical chips in U, and as the seeing degraded, we could not get a guidestar to lock on, even though we could see one that we thought should work near the edge of the chip. The seeing degradation as the wind rose probably makes this a mis-observation and I have provisionally left it in the queue (it requires only 30 minutes for each of the two OBs) as "not done", but this should be reviewed.

  • Standard star fields were observed only at the start, the mid-night field was aborted due to high winds.

  • Got sky flats at UBV at the start of the night, end-of-night flats were impossible due to high wind restrictions.

Equipment and the telescope performed well, despite problems as noted in the log above. PartnerObserving efficiency was pretty good, given this was my first for-real night with the system.

-- RichardPogge - 23 Jan 2007

Topic revision: r10 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback